# Improved Constructions of Reed-Solomon Codes with Optimal Repair Bandwidth

### v1

@@ Appendix B Proof of Lemma 3
$K=\sum_{u=0}^{s-1}\alpha^{u}S=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(u,v):(u,v)\in\mathcal{I}(s,p)\}.$
From previous analysis, we only need to prove that for $(i,j)\in\mathcal{I}(s,b_{1})$, $R(i,a_{0}s+j)\in K$.
Fix $(i,j)\in\mathcal{I}(s,b_{1})$, let $\tau(\cdot,\cdot)$ be as in Lemma 4, then
$R(i,j):=\alpha^{a_{0}s+i+j}\beta^{\tau(i,a_{0}s+j)}.$
Equivalently, if $(i,j)$ lies in the unique Euclidean square whose upper-left corner is $(y,x)$, then Lemma 4 gives $\tau(i,a_{0}s+j)=y+(j-x)=j+(y-x)$, hence
$R(i,j)=\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{\,y-x}.$
Since $\mathcal{R}_{j}\in S$, for every $i\in[0,s-1]$ we have $\alpha^{i}\mathcal{R}_{j}\in K$. By Proposition 1 together with Lemma  4, the element $\alpha^{i}\mathcal{R}_{j}$ has the form
$\alpha^{i}\mathcal{R}_{j}=\alpha^{a_{0}s+i}(\alpha\beta)^{j}\Big(\beta^{y-x}+\sum_{\begin{subarray}{c}Q\text{ intersects col }j\\ Q\neq Q_{0}\end{subarray}}\beta^{y_{Q}-x_{Q}}\Big),$
where $Q_{0}$ is the Euclidean square containing $(i,j)$ and, for each other intersected square $Q$, $(y_{Q},x_{Q})$ denotes its upper-left corner. Consequently,
$R(i,j)=\alpha^{i}\mathcal{R}_{j}-\alpha^{a_{0}s+i}(\alpha\beta)^{j}\sum_{\begin{subarray}{c}Q\text{ intersects col }j\\ Q\neq Q_{0}\end{subarray}}\beta^{y_{Q}-x_{Q}}.$ (15)
Therefore it is enough to show that for every $Q\neq Q_{0}$ intersecting column $j$,
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}\in K.$ (16)
We prove (16) by induction on the column index $j$ ordered by the sequence of quotients appearing in the Euclidean recursion. Let $j^{(0)}:=j$ and $j^{(2t)}$ be defined by (13). Order columns by the lexicographic order of the finite sequence
$\Big(\Big\lfloor\frac{j^{(0)}}{b_{2}}\Big\rfloor,\ \Big\lfloor\frac{j^{(2)}}{b_{4}}\Big\rfloor,\ \Big\lfloor\frac{j^{(4)}}{b_{6}}\Big\rfloor,\ \ldots\Big),$
where the sequence stops when the reachability condition 14 first fails. Assume that for every column $j^{\prime}$ preceding $j$ in this order and every $i^{\prime}\in[0,s-1]$, we already have $R(i^{\prime},j^{\prime})\in K$.
Fix an intersected square $Q\neq Q_{0}$ in column $j$. If $Q$ is a terminating square arising from a horizontal decomposition step, then its upper-left column coordinate is of the form $x_{Q}\equiv 0\pmod{b_{2t}}$ for the corresponding step size $b_{2t}$, and the contribution of $Q$ in $\alpha^{i}\mathcal{R}_{j}$ is the term $\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}$ (this is exactly the terminating term in (1)). Choose an integer $z\geq 1$ such that $(i+zb_{2t},\,j-zb_{2t})\in\mathcal{I}(s,b_{1})$ lies in the same terminating square pattern at that step (such a $z$ exists because the terminating square has side length $b_{2t}$ and column $j$ lies in a block with $\lfloor j^{(2t-2)}/b_{2t}\rfloor>0$ at that step). Then, by the definition of $r(\cdot,\cdot)$ via Lemma 4 and the identity
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=\alpha^{a_{0}s+i+zb_{2t}}(\alpha\beta)^{j-zb_{2t}}\beta^{y_{Q}-(x_{Q}-zb_{2t})}=\alpha^{a_{0}s+(i+zb_{2t})+(j-zb_{2t})}\beta^{\tau(i+zb_{2t},a_{0}s+j-zb_{2t})},$
we obtain
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=R(i+zb_{2t},\,j-zb_{2t}).$
Moreover, $j-zb_{2t}$ precedes $j$ in the above lexicographic order because the quotient $\lfloor j^{(2t-2)}/b_{2t}\rfloor$ strictly decreases while all earlier quotients remain unchanged. Hence, by the induction hypothesis, $R(i+zb_{2t},\,j-zb_{2t})\in K$, which proves (16) for terminating squares.
If $Q$ is a vertical square in the Euclidean partition, then its contribution in (1) belongs to one of the vertical progressions at some vertical step, and the corresponding exponent difference $y_{Q}-x_{Q}$ differs from the target difference yxy-x by a positive multiple of the vertical step size at that stage. In this case, the element $\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}$ can be rewritten as
$\alpha^{a_{0}s+i}(\alpha\beta)^{j}\beta^{y_{Q}-x_{Q}}=(\alpha\beta)^{j+(y_{Q}-x_{Q})}\alpha^{a_{0}s+i-(y_{Q}-x_{Q})}.$
Since $Q\neq Q_{0}$ and $(i,j)\in Q_{0}$, the row index $i$ lies strictly above the vertical block indexed by $Q$ at that step, hence $0\leq a_{0}s+i-(y_{Q}-x_{Q})\leq a_{0}s-1$. Therefore $(\alpha\beta)^{j+(y_{Q}-x_{Q})}\alpha^{a_{0}s+i-(y_{Q}-x_{Q})}\in K$ by $\sum_{i=0}^{s-1}\alpha^{i}S^{(1)}\in K$, proving (16) for vertical squares.
Thus every term in the sum on the right-hand side of (15) belongs to $K$, so (15) implies $R(i,j)\in K$. By induction over the above column order, we have $R(i,j)\in K$ for all $(i,j)\in\mathcal{I}(s,b_{1})$, which finished the proof.

### v2

@@ B-C Two algebraic facts on K
$\operatorname{Span}_{\mathbb{F}}\{\alpha^{z}\beta^{\theta},\dots,\alpha^{z+p-1}\beta^{\theta}\}=\mathbb{E}\beta^{\theta}\subseteq K.$
B-D Classification of Exp(j)
From Proposition 1, each $t\in\mathsf{Exp}(j)$ has exactly one of the following forms.
  1. 1.
    Vertical contribution with index $(t_{0},q)$. There exist $t_{0}\geq 1$ and $q\in[0,a_{2t_{0}-1}-1]$ such that $y_{2k}(j)=a_{2k}$ for all $k\in\{1,\dots,t_{0}-1\}$ and
    $t=\underbrace{\sum_{k=1}^{t_{0}-1}a_{2k-1}b_{2k-1}}_{=:Y_{t_{0}}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)}.$ (24)
  2. 2.
    Horizontal terminating contribution with index $t_{0}$. There exists $t_{0}\geq 1$ with $y_{2k}(j)=a_{2k}$ for all $k\in\{1,\dots,t_{0}-1\}$ and $y_{2t_{0}}(j)<a_{2t_{0}}$, such that
    $t=\underbrace{\sum_{k=1}^{t_{0}}a_{2k-1}b_{2k-1}}_{=:L_{t_{0}}}+j^{(2t_{0})}.$ (25)
Lemma 5 (Horizontal exponent is maximal in $\mathsf{Exp}(j)$).
For every $(i,j)\in\mathcal{I}(s,b_{1})$, let $t^{\star}:=\tau(i,a_{0}s+j)$. Then
[…]
Proof.
This is a direct corollary of the closed form of $\mathsf{Exp}(j)$ in Proposition 1.
B-E Definition of a locating function
Fix $\theta\in[0,s-1]$ and $k\in[0,b_{1}-1]$. Let
$Q_{\theta,k}=[y_{\theta,k},\,y_{\theta,k}+\lambda_{\theta,k}-1]\times[x_{\theta,k},\,x_{\theta,k}+\lambda_{\theta,k}-1]$
be the unique square in $\mathcal{P}(s,b_{1})$ containing $(\theta,k)$. Define
$\Phi_{\theta}(k):=(i_{\theta,k},j_{\theta,k}):=\bigl(y_{\theta,k}+(k-x_{\theta,k}),\;x_{\theta,k}+(\theta-y_{\theta,k})\bigr).$ (26)
Lemma 6 (Predecessors locating function $\Phi_{\theta}$).
For every $\theta\in[0,s-1]$ and $k\in[0,b_{1}-1]$,
$\tau(i_{\theta,k},a_{0}s+j_{\theta,k})=\theta,\qquad r(i_{\theta,k},j_{\theta,k})=\alpha^{a_{0}s+\theta+k}\beta^{\theta}.$
Equivalently,
$r\bigl(\Phi_{\theta}(k)\bigr)=\alpha^{a_{0}s+\theta+k}\beta^{\theta}.$
Proof.
Because $(\theta,k)\in Q_{\theta,k}$, there exist unique integers
$a:=\theta-y_{\theta,k},\qquad b:=k-x_{\theta,k},$
with $0\leq a,b\leq\lambda_{\theta,k}-1$. Hence $\Phi_{\theta}(k)=(y_{\theta,k}+b,\;x_{\theta,k}+a)\in Q_{\theta,k}$. Let
$Q=[y,y+\lambda-1]\times[x,x+\lambda-1]\in\mathcal{P}(s,b_{1}).$
In global column coordinates, the corresponding square of $R$ is
$Q^{\prime}=[y,y+\lambda-1]\times[a_{0}s+x,\;a_{0}s+x+\lambda-1].$
For $(u,v)\in Q$, write $u=y+r$, $v=x+c$ with $0\leq r,c\leq\lambda-1$. Then $(u,a_{0}s+v)\in Q^{\prime}$, so by Lemma 4 (applied on $Q^{\prime}$),
$\tau(u,a_{0}s+v)=y+c=y+(v-x).$
Applying this with $Q=Q_{\theta,k}$ and $(u,v)=\Phi_{\theta}(k)$ gives
$\tau(i_{\theta,k},a_{0}s+j_{\theta,k})=y_{\theta,k}+\big(j_{\theta,k}-x_{\theta,k}\big)=y_{\theta,k}+a=\theta.$
Therefore
$r(i_{\theta,k},j_{\theta,k})=\alpha^{a_{0}s+i_{\theta,k}+j_{\theta,k}}\beta^{\theta}.$
Finally,
$i_{\theta,k}+j_{\theta,k}=\big(y_{\theta,k}+b\big)+\big(x_{\theta,k}+a\big)=(a+b)+(x_{\theta,k}+y_{\theta,k})=\theta+k,$
so $r(i_{\theta,k},j_{\theta,k})=\alpha^{a_{0}s+\theta+k}\beta^{\theta}$. ∎
B-F Lemmas for locating predecessors (C1/C3)
Lemma 7 (Locating predecessors for Case C1).
Fix $(i,j)\in[0,s-1]\times[0,b_{1}-1]$ and let
$t\in\mathsf{Exp}(j)\setminus\{\tau(i,a_{0}s+j)\},\qquad t<t^{\star}:=\tau(i,a_{0}s+j).$
Then $t$ is vertical (by Lemma 5). Equivalently, there exist $t_{0}\geq 1$ and $q\in[0,a_{2t_{0}-1}-1]$ such that
$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)}.$
Moreover, $q<x_{2t_{0}-1}(i)$. Then, for every $k\in[0,b_{1}-1]$, the point $\Phi_{t}(k)=(i_{k},j_{k})$ satisfies
$(i_{k},j_{k})\prec(i,j).$
Proof.
By Lemma 5, $t$ is vertical, hence
$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)}.$
We now prove $q<x_{2t_{0}-1}(i)$.
If $x_{2t_{0}-1}(i)\leq a_{2t_{0}-1}-1$, then
$t^{\star}=\tau(i,a_{0}s+j)=Y_{t_{0}}+x_{2t_{0}-1}(i)\,b_{2t_{0}-1}+j^{(2t_{0}-2)}.$
Comparing with
$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)}$
and using $t<t^{\star}$, we get $q<x_{2t_{0}-1}(i)$.
If $x_{2t_{0}-1}(i)=a_{2t_{0}-1}$, then
$q\leq a_{2t_{0}-1}-1<x_{2t_{0}-1}(i),$
so again $q<x_{2t_{0}-1}(i)$.
Therefore
$q<x_{2t_{0}-1}(i).$ (27)
Set
$C:=\sum_{h=1}^{t_{0}-1}a_{2h}b_{2h}=b_{1}-b_{2t_{0}-1}.$
From $y_{2h}(j)=a_{2h}$ for $h<t_{0}$, we have
$j=C+j^{(2t_{0}-2)}\geq C.$
Case I: $0\leq k\leq C-1$. Write $\Phi_{t}(k)=(i_{k},j_{k})$ and let
$Q_{t,k}=[y_{t,k},y_{t,k}+\lambda_{t,k}-1]\times[x_{t,k},x_{t,k}+\lambda_{t,k}-1]$
be the unique square containing $(t,k)$. Since $k<C$ and $C=b_{1}-b_{2t_{0}-1}$, the square $Q_{t,k}$ is contained in $0,s-1]\times[0,C-1]$. Indeed, $x=C$ is a square-edge boundary after horizontal step $(2t_{0}-2)$; all later decomposition steps act only inside the residual strip $C,b_{1}-1]$. Since $k<C$, the point $(t,k)$ lies strictly to the left of this boundary, so its containing square cannot cross $x=C$. Hence
$x_{t,k}+\lambda_{t,k}\leq C.$
By (26),
$j_{k}=x_{t,k}+\bigl(t-y_{t,k}\bigr)\leq x_{t,k}+\lambda_{t,k}-1\leq C-1<j,$
where the first inequality uses $(t,k)\in Q_{t,k}$, which implies $0\leq t-y_{t,k}\leq\lambda_{t,k}-1$. Hence $j_{k}<j$, so by the column-priority consequence of $\prec$,
$(i_{k},j_{k})\prec(i,j).$
Case II: $C\leq k\leq b_{1}-1$. From
$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)},\qquad 0\leq j^{(2t_{0}-2)}\leq b_{2t_{0}-1}-1,$
we get
$Y_{t_{0}}+q\,b_{2t_{0}-1}\leq t\leq Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1.$
Also, $C\leq k\leq b_{1}-1$ and $C=b_{1}-b_{2t_{0}-1}$ imply
$C\leq k\leq C+b_{2t_{0}-1}-1.$
Hence
$(t,k)\in Q^{\ast},$
where
$Q^{\ast}=[Y_{t_{0}}+q\,b_{2t_{0}-1},\,Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1]\times[C,\,b_{1}-1].$
Hence $x_{t,k}=C$, $y_{t,k}=Y_{t_{0}}+q\,b_{2t_{0}-1}$, $\lambda_{t,k}=b_{2t_{0}-1}$ and
$t-y_{t,k}=j^{(2t_{0}-2)}.$
Therefore
$j_{k}=x_{t,k}+\bigl(t-y_{t,k}\bigr)=C+j^{(2t_{0}-2)}=j.$
Also
$i_{k}=y_{t,k}+(k-C)\in[Y_{t_{0}}+q\,b_{2t_{0}-1},\,Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1],$
which implies
$x_{2t_{0}-1}(i_{k})=q.$
Combining this with (27), we obtain
$x_{2t_{0}-1}(i_{k})<x_{2t_{0}-1}(i).$
For odd indices $1,3,\dots,2t_{0}-3$, the quotients of $i_{k}$ and $i$ coincide; hence
$\mathbf{X}(i_{k})<_{\rm lex}\mathbf{X}(i).$
Since $j_{k}=j$, this gives $(i_{k},j_{k})\prec(i,j)$.
Combining Case I and Case II yields $(i_{k},j_{k})\prec(i,j)$ for all $k\in[0,b_{1}-1]$. ∎
Lemma 8 (Locating a predecessor for Case C3).
Let $(i,j)\in[0,s-1]\times[0,b_{1}-1]$, let
$t^{\star}:=\tau(i,a_{0}s+j),\qquad t\in\mathsf{Exp}(j)\setminus\{t^{\star}\},$
and assume
$t>t^{\star},\qquad i+j\geq t.$
Define
$\delta:=i+j-t.$
Then:
  1. (a)
    $\delta\in[0,b_{1}-1]$.
  2. (b)
    The point $(i_{\delta},j_{\delta}):=\Phi_{t}(\delta)$ satisfies
    $r(i_{\delta},j_{\delta})=\alpha^{a_{0}s+i+j}\beta^{t},\qquad(i_{\delta},j_{\delta})\prec(i,j).$
Proof.
Proof of (a). Since $i+j\geq t$, we have $\delta\geq 0$. It remains to show $\delta\leq b_{1}-1$.
Case A1: $t$ is vertical. Write
$t=Y_{t_{0}}+q\,b_{2t_{0}-1}+j^{(2t_{0}-2)},\qquad C:=\sum_{h=1}^{t_{0}-1}a_{2h}b_{2h}=b_{1}-b_{2t_{0}-1},\qquad j=C+j^{(2t_{0}-2)}.$
Let $Q_{t,j}$ be the unique square in $\mathcal{P}(s,b_{1})$ containing $(t,j)$. From $0\leq j^{(2t_{0}-2)}\leq b_{2t_{0}-1}-1$, we have
$Y_{t_{0}}+q\,b_{2t_{0}-1}\leq t\leq Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1,\qquad C\leq j\leq b_{1}-1.$
Hence $(t,j)$ lies in
$Q_{v}:=[Y_{t_{0}}+q\,b_{2t_{0}-1},\,Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1]\times[C,b_{1}-1].$
At vertical step $(2t_{0}-1)$, $Q_{v}$ is one square of $\mathcal{P}(s,b_{1})$; therefore $Q_{t,j}=Q_{v}$. By Lemma 4, for all
$u\in[Y_{t_{0}}+q\,b_{2t_{0}-1},\,Y_{t_{0}}+(q+1)b_{2t_{0}-1}-1],$
we have
$\tau(u,a_{0}s+j)=Y_{t_{0}}+q\,b_{2t_{0}-1}+(j-C)=t.$
For fixed $j$, Proposition 1 orders exponents by Euclidean terms, so rows below this interval satisfy $\tau(u,a_{0}s+j)>t$. Since $t^{\star}=\tau(i,a_{0}s+j)<t$, row $i$ cannot be in this interval or below it. Hence
$i\leq Y_{t_{0}}+q\,b_{2t_{0}-1}-1.$
Therefore
$\delta=i+j-t=i+C-\bigl(Y_{t_{0}}+q\,b_{2t_{0}-1}\bigr)\leq C-1\leq b_{1}-1.$
Case A2: $t$ is horizontal. Write
$t=L_{t_{0}}+j^{(2t_{0})},$
where
$C_{h}:=\sum_{h=1}^{t_{0}-1}a_{2h}b_{2h}+y_{2t_{0}}(j)b_{2t_{0}}.$
By the remainder decomposition,
$j=\sum_{h=1}^{t_{0}-1}a_{2h}b_{2h}+j^{(2t_{0}-2)},\qquad j^{(2t_{0}-2)}=y_{2t_{0}}(j)b_{2t_{0}}+j^{(2t_{0})}.$
Hence
$j=C_{h}+j^{(2t_{0})}.$
Let $Q_{t,j}$ be the unique square in $\mathcal{P}(s,b_{1})$ containing $(t,j)$. From $0\leq j^{(2t_{0})}\leq b_{2t_{0}}-1$, we have $(t,j)\in Q_{h}$, where
$Q_{h}:=[L_{t_{0}},\,L_{t_{0}}+b_{2t_{0}}-1]\times[C_{h},\,C_{h}+b_{2t_{0}}-1].$
At horizontal step $(2t_{0})$, $Q_{h}$ is one square of $\mathcal{P}(s,b_{1})$; hence $Q_{t,j}=Q_{h}$. By Lemma 4, for all $u\in[L_{t_{0}},L_{t_{0}}+b_{2t_{0}}-1]$,
$\tau(u,a_{0}s+j)=L_{t_{0}}+(j-C_{h})=L_{t_{0}}+j^{(2t_{0})}=t.$
By Euclidean recursion,
$s=b_{0}=\sum_{h=1}^{t_{0}}a_{2h-1}b_{2h-1}+b_{2t_{0}}=L_{t_{0}}+b_{2t_{0}},$
so $L_{t_{0}},L_{t_{0}}+b_{2t_{0}}-1]=[L_{t_{0}},s-1]$. If $i\geq L_{t_{0}}$, then $\tau(i,a_{0}s+j)=t$, contradicting $t^{\star}<t$. Hence $i\leq L_{t_{0}}-1$, and
$\delta=i+j-t=i+C_{h}-L_{t_{0}}\leq C_{h}-1.$
Since $j=C_{h}+j^{(2t_{0})}$ and $0\leq j\leq b_{1}-1$, we have $C_{h}\leq b_{1}-1$, therefore
$\delta\leq b_{1}-1.$
Combining A1 and A2, we get $\delta\in[0,b_{1}-1]$.
Proof of (b). Set $(i_{\delta},j_{\delta}):=\Phi_{t}(\delta)$. By Lemma 6,
$r(i_{\delta},j_{\delta})=\alpha^{a_{0}s+t+\delta}\beta^{t}=\alpha^{a_{0}s+i+j}\beta^{t}.$
It remains to prove $(i_{\delta},j_{\delta})\prec(i,j)$.
Case B1: $t$ is vertical. From part (a), $\delta\leq C-1$. Let
$Q_{t,\delta}=[y_{t,\delta},y_{t,\delta}+\lambda_{t,\delta}-1]\times[x_{t,\delta},x_{t,\delta}+\lambda_{t,\delta}-1]$
be the unique square containing $(t,\delta)$. Since $(t,\delta)\in Q_{t,\delta}$,
$0\leq t-y_{t,\delta}\leq\lambda_{t,\delta}-1,\qquad 0\leq\delta-x_{t,\delta}\leq\lambda_{t,\delta}-1.$
The line $x=C$ is a square-edge boundary after step $2t_{0}-2$. Because $\delta<C$, $Q_{t,\delta}$ cannot cross this boundary, hence
$x_{t,\delta}+\lambda_{t,\delta}\leq C.$
Therefore
$j_{\delta}=x_{t,\delta}+(t-y_{t,\delta})\leq x_{t,\delta}+\lambda_{t,\delta}-1\leq C-1.$
Also $j=C+j^{(2t_{0}-2)}$, so
$j\geq C.$
Therefore $j_{\delta}\leq C-1<j$, and by the column-priority consequence of $\prec$,
$(i_{\delta},j_{\delta})\prec(i,j).$
Case B2: $t$ is horizontal. From part (a), $\delta\leq C_{h}-1$. Let
$Q_{t,\delta}=[y_{t,\delta},y_{t,\delta}+\lambda_{t,\delta}-1]\times[x_{t,\delta},x_{t,\delta}+\lambda_{t,\delta}-1]$
be the unique square containing $(t,\delta)$. Since $(t,\delta)\in Q_{t,\delta}$,
$0\leq t-y_{t,\delta}\leq\lambda_{t,\delta}-1,\qquad 0\leq\delta-x_{t,\delta}\leq\lambda_{t,\delta}-1.$
The line $x=C_{h}$ is a square-edge boundary at horizontal step $2t_{0}$. Because $\delta<C_{h}$, $Q_{t,\delta}$ cannot cross this boundary, so
$x_{t,\delta}+\lambda_{t,\delta}\leq C_{h}.$
Hence
$j_{\delta}=x_{t,\delta}+(t-y_{t,\delta})\leq C_{h}-1.$
Write $j=C_{h}+j^{(2t_{0})}$. Since $j\geq C_{h}$, we have $j_{\delta}\leq C_{h}-1<j$, hence by the column-priority consequence of $\prec$,
$(i_{\delta},j_{\delta})\prec(i,j).$
B-G Main induction
We prove (19) for all $(i,j)\in[0,s-1]\times[0,b_{1}-1]$ by induction on $\prec$.
Initialization.
For $j=0$ and $i\in[0,b_{1}-1]$, every interference exponent $t\in\mathsf{Exp}(0)\setminus\{0\}$ satisfies $t\geq b_{1}>i=i+j$. Hence
$\alpha^{a_{0}s+i}\beta^{t}=(\alpha\beta)^{t}\alpha^{a_{0}s+i-t}\in K$
by Fact 1, so $r(i,0)\in K$ on this initial boundary set.
Inductive step.
Fix $(i,j)$ and assume
$r(i^{\prime},j^{\prime})\in K\quad\forall(i^{\prime},j^{\prime})\prec(i,j).$
Set $t^{\star}:=\tau(i,a_{0}s+j)$, and take any $t\in\mathsf{Exp}(j)\setminus\{t^{\star}\}$.
We show
$\alpha^{a_{0}s+i+j}\beta^{t}\in K.$
Case C1: $t<t^{\star}$. By Lemma 5, this branch contains no horizontal term; thus $t$ is vertical and Lemma 7 applies.
First, Fact 1 gives
$\alpha^{t+u}\beta^{t}\in K,\qquad u=0,1,\dots,a_{0}s-1.$
Next, for each $k\in[0,b_{1}-1]$, let $(i_{k},j_{k})=\Phi_{t}(k)$. By Lemma 7, $(i_{k},j_{k})\prec(i,j)$, so by induction hypothesis,
$r(i_{k},j_{k})\in K.$
By Lemma 6,
$r(i_{k},j_{k})=\alpha^{a_{0}s+t+k}\beta^{t},$
thus
$\alpha^{a_{0}s+t+k}\beta^{t}\in K,\qquad k=0,1,\dots,b_{1}-1.$
Therefore we have $p=a_{0}s+b_{1}$ consecutive monomials
$\alpha^{t}\beta^{t},\alpha^{t+1}\beta^{t},\dots,\alpha^{t+p-1}\beta^{t}\in K.$
By Fact 2,
$\beta^{t}\mathbb{E}\subseteq K.$
In particular, $\alpha^{a_{0}s+i+j}\beta^{t}\in K$.
Case C2: $t>t^{\star}$ and $i+j<t$. Then
$\alpha^{a_{0}s+i+j}\beta^{t}=(\alpha\beta)^{t}\alpha^{a_{0}s+i+j-t},$
with $0\leq a_{0}s+i+j-t\leq a_{0}s-1$. Hence Fact 1 gives $\alpha^{a_{0}s+i+j}\beta^{t}\in K$.
Case C3: $t>t^{\star}$ and $i+j\geq t$. Let $\delta:=i+j-t$. By part (a) of Lemma 8, $\delta\in[0,b_{1}-1]$. By part (b) of Lemma 8, with $(i_{\delta},j_{\delta})=\Phi_{t}(\delta)$,
$r(i_{\delta},j_{\delta})=\alpha^{a_{0}s+i+j}\beta^{t},\qquad(i_{\delta},j_{\delta})\prec(i,j).$
So induction gives $r(i_{\delta},j_{\delta})\in K$, i.e.
$\alpha^{a_{0}s+i+j}\beta^{t}\in K.$
Thus in Cases C1, C2 and C3 every interference monomial in (21) belongs to $K$. Together with (20), this yields $r(i,j)\in K$. The induction is complete.
B-H Conclusion
By induction over $\prec$, (19) holds for all $(i,j)\in[0,s-1]\times[0,b_{1}-1]$. Hence every entry of $R$ belongs to $K$. Since $R$ is a rearrangement of entries of $B$, we have $\mathbb{K}=\operatorname{Span}_{\mathbb{F}}(B)\subseteq K$. As $K\subseteq\mathbb{K}$ by definition, $K=\mathbb{K}$, namely
$\mathbb{K}=\sum_{u=0}^{s-1}\alpha^{u}S.$
$\square$
B-I Concrete verification for $m=4$ (all $a_{i}=2$)
Take p=70,s=29p=70,\ s=29, so the Euclidean chain is
$70=2\cdot 29+12,\quad 29=2\cdot 12+5,\quad 12=2\cdot 5+2,\quad 5=2\cdot 2+1,\quad 2=2\cdot 1.$
Hence
$b_{1}=12,\ b_{2}=5,\ b_{3}=2,\ b_{4}=1,\qquad a_{0}s=58.$
Exponent array on $R_{1}$ (last $b_{1}$ columns of $R$). Since $a_{0}s=58$, these columns are indexed by global columns $58+j$ with $0\leq j\leq 11$. For $(i,j)\in[0,28]\times[0,11]$, we have
$\tau(i,58+j)=\begin{cases}j,&0\leq i\leq 11,\\ 12+j,&12\leq i\leq 23,\\ 24+(j\bmod 5),&24\leq i\leq 28,\ 0\leq j\leq 9,\\ 24+(j-10),&24\leq i\leq 25,\ j\in\{10,11\},\\ 26+(j-10),&26\leq i\leq 27,\ j\in\{10,11\},\\ 28,&i=28,\ j\in\{10,11\}.\end{cases}$
For column $j=10$ in $R_{1}$ (global column index $58+10=68$), we have
$j^{(0)}=10,\quad y_{2}(10)=\lfloor 10/5\rfloor=2,\quad j^{(2)}=0,\quad y_{4}(10)=0,\quad j^{(4)}=0.$
Therefore
$\mathsf{Exp}(10)=\{10,22,24,26,28\},$
where $28$ is the unique horizontal term, and
$\tau(i,68)=\begin{cases}10,&0\leq i\leq 11,\\ 22,&12\leq i\leq 23,\\ 24,&24\leq i\leq 25,\\ 26,&26\leq i\leq 27,\\ 28,&i=28.\end{cases}$
Equivalently,
$(\tau(0,68),\dots,\tau(28,68))=(\underbrace{10,\dots,10}_{12},\underbrace{22,\dots,22}_{12},24,24,26,26,28).$
Five-segment repair on column $j=10$ of $R_{1}$. Let
$I_{1}=[0,11],\quad I_{2}=[12,23],\quad I_{3}=[24,25],\quad I_{4}=[26,27],\quad I_{5}=\{28\}.$
On these segments,
$t^{\star}(i):=\tau(i,68)=10,22,24,26,28\quad\text{for }i\in I_{1},I_{2},I_{3},I_{4},I_{5}.$
For each $i\in[0,28]$ and each $t\in\mathsf{Exp}(10)\setminus\{t^{\star}(i)\}$, we split by Cases C1, C2, C3.
Segment $I_{1}$ ($t^{\star}=10$). For $t\in\{22,24,26,28\}$ and $i\in I_{1}$,
$i+10\leq 21<t,$
so this segment is entirely Case C2.
Segment $I_{2}$ ($t^{\star}=22$). For $t=10$, this is Case C1. For $t\in\{24,26,28\}$, it is Case C2 when $i+10<t$, and Case C3 when $i+10\geq t$, i.e.
[…]
Segment $I_{3}$ ($t^{\star}=24$). $t\in\{10,22\}$ gives Case C1; $t\in\{26,28\}$ gives Case C3, since $i+10\geq 34$.
Segment $I_{4}$ ($t^{\star}=26$). $t\in\{10,22,24\}$ gives Case C1; $t=28$ gives Case C3.
Segment $I_{5}$ ($t^{\star}=28$). All $t\in\{10,22,24,26\}$ are Case C1.
Thus every interference term in column $j=10$ is covered by one of C1/C2/C3. Case C2 is handled by Fact 1, Case C1 by Lemma 7, and Case C3 by Lemma 8.
How one concrete entry is recovered: $(i,j)=(26,10)$. Here
$t^{\star}=\tau(26,68)=26,\qquad\mathsf{Exp}(10)\setminus\{t^{\star}\}=\{10,22,24,28\}.$
Since $26+10=36>28$, there is no C2 term; $t=10,22,24$ are C1 and $t=28$ is C3. We show explicitly that the C1 representative $t=24$ and the C3 term $t=28$ are both eliminable.
C1 representative ($t=24$). For $k\in[0,11]$, let $(i_{k},j_{k})=\Phi_{24}(k)$. For example, take $k=8$. In this instance, $(24,8)$ lies in the square
$Q_{24,8}=[24,28]\times[5,9],$
so $y_{24,8}=24$, $x_{24,8}=5$, and by (26),
$\Phi_{24}(8)=\big(24+(8-5),\;5+(24-24)\big)=(27,5)\prec(26,10).$
In fact, by Lemma 7,
$(i_{k},j_{k})\prec(26,10)\quad\text{for all }k.$
By Lemma 6,
$r(i_{k},j_{k})=\alpha^{58+24+k}\beta^{24},$
so $\alpha^{58+24}\beta^{24},\dots,\alpha^{58+35}\beta^{24}\in K$. Together with the initial block from Fact 1, Fact 2 implies that $\beta^{24}\mathbb{E}\subseteq K$, hence
$\alpha^{58+26+10}\beta^{24}\in K.$
Therefore, the interference term with exponent $t=24$ at $(26,10)$ can be canceled.
C3 term ($t=28$). Set $\delta=26+10-28=8$. By Lemma 8, part (b), with $(i_{\delta},j_{\delta})=\Phi_{28}(8)$,
$r(i_{\delta},j_{\delta})=\alpha^{58+26+10}\beta^{28},\qquad(i_{\delta},j_{\delta})\prec(26,10).$
To compute this point explicitly: $(28,8)$ is in
$Q_{28,8}=[24,28]\times[5,9],$
hence $y_{28,8}=24$, $x_{28,8}=5$, and by (26),
$\Phi_{28}(8)=\big(24+(8-5),\;5+(28-24)\big)=(27,9).$
Thus its column index is $9<10$, so this predecessor is already recovered. Hence
$\alpha^{58+26+10}\beta^{28}\in K,$
which cancels the $t=28$ interference at $(26,10)$.
The remaining terms $t=10,22$ are also C1 and are handled identically to $t=24$. So every interference term at the specific entry $(26,10)$ is removable, exactly as required by the induction proof.